If a train 1 kilometer long is passing through a tunnel 7 kilometers long at a speed of 80 kilometers per hour how long will it take the train to pass completely through the tunnel?
Answer by mananth(16949)

Train is 1 km
Tunnel is 7 km
The end of the train has to be out of the tunnel
Distance = 7+1 = 8km
speed = 80 km/h
time = distance / speed
time = 8/80
time = 1/8 of an hour
=7.5 minutes